OpenStudy (anonymous):

solve using the addition principle -42+y=3

OpenStudy (anonymous):

y = 3 + 42 y = 45

OpenStudy (anonymous):

add 42 to each side. ex. 1. Solve: x + 6 = -15 Solution: Using the Addition Principle, add -6 to each side of the equation. x + 6 - 6 = -15 - 6 The variable is now isolated. x = -21

OpenStudy (anonymous):

you can also call it the additive inverse